A poster in the shape of a triangle has one side that is five inches more the length of the shortest side, and another side that is three inches less than twice the shortest side. Find the dimensions of the poster if its perimeter is 42 inches
Answer by Paul(988)

The shortest side is x
The oter two sides are, x+5, and 2x-3
Equation:
(x)+(x+5)+(2x-3)=42
4x+2=42
4x=40
x=10
10+5=15
2(10)-3=17
Hence, the three sides are 10, 15 and 17 cm.
Paul.